OpenStudy (anonymous):

Solve by the linear combination method (with or without multiplication). 2x + 3y = 11 –4x + 2y = 2 A. (1, 3) B. (2, 3) C. (1, 2) D. (1, –2)

OpenStudy (anonymous):

I think it is A, Right?

OpenStudy (anonymous):

2x + 3y = 11 2( 2x + 3y = 11) 4x + 6y = 22 –4x + 2y = 2 -4x + 2y = 2 8y = 24 y = 3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thank you can you answer another one for me?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Solve by the linear combination method (with or without multiplication). x + y = 40 0.08x + 0.03y = 1.7 A. (40, 60) B. (20, 20) C. (10, 30) D. (20, –50)

OpenStudy (anonymous):

What do you think it is?

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

y = 40-x 0.08x + 0.03 (40-x) = 1.7 solve this...it's not B

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

I got C.

OpenStudy (anonymous):

0.08x + 1.2 - 0.3x = 1.7 0.05x = 1.5 x = 10 yes...C

OpenStudy (anonymous):

ok one more...please?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Solve by the linear combination method (with or without multiplication). 3x – 4y = 37 5x + 2y = 1 A. (9, 4) B. (3, 7) C. (–3, –7) D. (3, –7)

OpenStudy (anonymous):

Show me what you're trying

OpenStudy (anonymous):

I got D.

OpenStudy (anonymous):

Me, too. Good

OpenStudy (anonymous):

thanks again!